$SPAD/src/input radix.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool radix.output
)set message test on
)set message auto off
)clear all
--S 1 of 17
111::RadixExpansion(5)
--R
--R
--R   (1)  421
--R                                                      Type: RadixExpansion(5)
--E 1
--S 2 of 17
(5/24)::RadixExpansion(2)
--R
--R
--R             __
--R   (2)  0.00110
--R                                                      Type: RadixExpansion(2)
--E 2
--S 3 of 17
(5/24)::RadixExpansion(3)
--R
--R
--R           __
--R   (3)  0.012
--R                                                      Type: RadixExpansion(3)
--E 3
--S 4 of 17
(5/24)::RadixExpansion(8)
--R
--R
--R           __
--R   (4)  0.152
--R                                                      Type: RadixExpansion(8)
--E 4

--S 5 of 17
(5/24)::RadixExpansion(10)
--R
--R
--R             _
--R   (5)  0.2083
--R                                                     Type: RadixExpansion(10)
--E 5
--S 6 of 17
(5/24)::RadixExpansion(12)
--R
--R
--R   (6)  0.26
--R                                                     Type: RadixExpansion(12)
--E 6
--S 7 of 17
(5/24)::RadixExpansion(16)
--R
--R
--R           _
--R   (7)  0.35
--R                                                     Type: RadixExpansion(16)
--E 7
--S 8 of 17
(5/24)::RadixExpansion(36)
--R
--R
--R   (8)  0.7I
--R                                                     Type: RadixExpansion(36)
--E 8
--S 9 of 17
(5/24)::RadixExpansion(38)
--R
--R
--R                    _____
--R   (9)  0 . 7 34 31 25 12
--R                                                     Type: RadixExpansion(38)
--E 9
--S 10 of 17
a := (76543/210)::RadixExpansion(8)
--R
--R
--R              ____
--R   (10)  554.37307
--R                                                      Type: RadixExpansion(8)

--E 10
--S 11 of 17
w := wholeRagits a
--R
--R
--R   (11)  [5,5,4]
--R                                                          Type: List(Integer)
--E 11
--S 12 of 17
f0 := prefixRagits a
--R
--R
--R   (12)  [3]
--R                                                          Type: List(Integer)
--E 12
--S 13 of 17
f1 := cycleRagits a
--R
--R
--R   (13)  [7,3,0,7]
--R                                                          Type: List(Integer)
--E 13
--S 14 of 17
u:RadixExpansion(8):=wholeRadix(w)+fractRadix(f0,f1)
--R
--R
--R              ____
--R   (14)  554.37307
--R                                                      Type: RadixExpansion(8)
--E 14
--S 15 of 17
v: RadixExpansion(12) := fractRadix([1,2,3,11], [0])
--R
--R
--R               _
--R   (15)  0.123B0
--R                                                     Type: RadixExpansion(12)
--E 15
--S 16 of 17
fractRagits(u)
--R
--R
--R              _______
--R   (16)  [3,7,3,0,7,7]

--R                                                        Type: Stream(Integer)
--E 16
--S 17 of 17
a :: Fraction(Integer)
--R
--R
--R         76543
--R   (17)  -----
--R          210
--R                                                      Type: Fraction(Integer)
--E 17
)spool
)lisp (bye)

———-

References^{[1] nothing}

7